OpenStudy (anonymous):

3 log x - 1/4 log (x^2+1) + 4 log (x-1) COMBINE

OpenStudy (anonymous):

\[log(x ^{3}-\log(x ^{2}+1)^{1/4}+\log(x-1)^{^{4}}\]\[\log(x ^{2}+1)^{1/4}/x ^{3}+\log (x-1)^{4}\] \[\log[((x ^{2} +1)^{1/4}/x ^{3})\times(x-1)^{4}]\] this is my idea. may be it's not right.